OpenStudy (anonymous):

Factor by grouping 3rs - 2r - 3s + 2

OpenStudy (callisto):

3rs - 2r - 3s + 2 = r (3r -2) - (3r - 2) = ... ?

OpenStudy (anonymous):

(3rs-2r)-(3s-2) r(3s-2)-(3s-2) (r-1)(3s-2)